# Geometry-Controlled Work Extraction in a Non-Markovian Quantum Battery

### v1

@@ Abstract
Introduction.— In recent years the thermodynamic effects and consequences arising from quantum evolution have emerged as a fundamental framework for the development of novel architectures for high-efficiency energy storage devices. These advances also provide the basis for the formulation and optimization of precise protocols for the control, conversion, and management of work and heat, approaches that open perspectives beyond the intrinsic limitations of classical thermodynamic processes [1, 2, 3, 4].
@@
Consequently, they have been extensively investigated and systematically analyzed in the contemporary literature [4, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Quantum batteries have been theoretically and experimentally investigated across a broad range of physical realizations. Proposed schemes include spin-chain models [15, 16, 17, 18] , atomic ensembles coupled to optical cavities [19, 20, 21, 22, 23] , and superconducting systems [24, 25, 26] . These diverse approaches rely on different interaction mechanisms and control techniques, thereby enabling the exploration of multiple strategies for microscopic energy storage and controlled energy transfer within quantum regimes.
@@
In this work, we investigate the role of the relative geometry between the charger and the quantum battery in a structured environment. In particular, we examine how the spatial seperation between the two qubits and the associated geometric phase $k_{0}d$ influence the energy transfer process, charging dynamics and the amount of the extractable work of the battery through their control over collective interference and the spectral structure of the environment. Within this framework, we demonstrate that the system geometry not only regulates the effective coupling strength between the charger and the environment, but also, through the engineering of bright and dark collective channels, can suppress or enhance decay processes and energy backflow. As a result, the thermodynamic performance of the battery, including charging time, charging power, and maximum ergotropy, becomes directly governed by the geometric distance parameter. Furthermore, our analysis reveals that in the presence of environmental memory (non-Markovian effects), the system geometry can give rise to distinct regions in parameter space where the revival of extractable work occurs. In this sense, geometry emerges as a fundamental control resource, enabling the active engineering and optimization of quantum battery performance.
Figures of merit..— In this section, we present the thermodynamic framework employed to analyze quantum batteries. The system under consideration is a quantum battery characterized by the Hamiltonian $H_{B}$, with its state described by the density operator $\rho_{B}$. This quantum state can be expressed through its spectral decomposition in terms of its eigenvalues $r_{j}$ and corresponding eigenvectors $|r_{j}\rangle$
@@
Physical model.— We consider a composite quantum system consisting of two identical two-level systems (TLSs) positioned at $x_{1}=-d$ and $x_{2}=+d$, respectively. Within the quantum battery paradigm, one TLS is designated as the charger, while the other functions as the battery. This physical setup is equivalent to two qubits placed inside a mirror-terminated one-dimensional waveguide of length $L$, where the boundary conditions discretize the electromagnetic field into a uniformly spaced ladder of standing-wave modes characterized by $k_{n}=n\pi/L$. The mode spacing is $\delta\omega=\pi v/L$, with $v$ denoting the group velocity. Such discretization introduces a natural environmental timescale associated with field rephasing, which plays a crucial role in the energy backflow and work revival phenomena observed in the quantum battery. We set $\hbar=1$ and work within the dipole and rotating-wave approximations. The total Hamiltonian of the system can be written as
@@
Figure 1: Schematic of the physical model: two identical two-level systems separated by a distance $2d$ inside a mirror-terminated one-dimensional waveguide. Qubit 1 acts as the charger (initially excited) and qubit 2 as the quantum battery (initially in the ground state). The spatial phase $k_{0}d$ controls collective interference and energy transfer dynamics.
where $H_{S}$ describes the free Hamiltonian of the two qubits (charger and battery), $H_{B}$ represents the discrete bosonic modes of the structured waveguide environment, and $H_{I}$ accounts for the interaction between the qubits and the environmental modes. The system Hamiltonian reads [39]
$H_{S}=\omega_{0}(\sigma_{1}^{+}\sigma_{1}^{-}+\sigma_{2}^{+}\sigma_{2}^{-})+J(\sigma_{1}^{+}\sigma_{2}^{-}+\sigma_{1}^{-}\sigma_{2}^{+}),$ (7)
where $\omega_{0}$ is the transition frequency of the two-level systems. The parameter $J$ represents the coherent dipole-dipole interaction [40]. In the quantum battery interpretation adopted here, qubit $1$ acts as the charger, while qubit $2$ represents the battery. The operators $\sigma_{j}^{\pm}$ denote the raising and lowering operators of the $j$-th qubit. The bath Hamiltonian is given by
@@
$H_{SB}=\sum_{k}\left[g_{k}^{(1)}\sigma_{1}^{-}a_{k}^{\dagger}+g_{k}^{(2)}\sigma_{2}^{-}a_{k}^{\dagger}+h.c.\right].$ (9)
@@
The model is schematically illustrated in Fig. 1 The influence of geometry is encoded in the coupling amplitudes $g_{k}^{(j)}=g_{k}e^{ikx_{j}}$, with $x_{j}$ representing the location of the $j$-th qubit [41, 42, 43, 44] . The resulting phase difference between the qubits generates interference in the collective channels, enabling controlled modulation of both charging dynamics and work extraction. Since the interaction of the qubits predominantly occurs with modes near their transition frequency $\omega_{0}$, we simplify the analytical treatment by approximating the wave number in the position-dependent phase terms as $k\approx k_{0}=\omega_{0}/v=2\pi/\lambda_{0}$, while retaining the full dispersion relation $\omega_{k}$ elsewhere in the Hamiltonian. Here, $\lambda_{0}$ denotes the resonant wavelength associated with the qubit transition frequency. Consequently, the position-dependent coupling coefficients can be expressed as
$g_{k}^{(1)}\simeq g_{k}e^{+ik_{0}d},\quad g_{k}^{(2)}\simeq g_{k}e^{-ik_{0}d}.$ (10)
In this form, the geometric dependence of the system is explicitly encoded in the phase factor $k_{0}d$. This phase plays a crucial role in controlling the interference between collective interaction channels and, in turn, determines the energy transfer dynamics and the amount of extractable work in the quantum battery. This approximation introduces a negligible error for physically relevant parameter regimes, remaining well below typical numerical thresholds. Consequently, it can be regarded as a well-controlled and reliable procedure. Widely adopted in studies of collective emission and waveguide quantum electrodynamics, this substitution preserves all geometry-induced interference effects while substantially simplifying the mathematical treatment. As a result, it enables closed-form analytical solutions in both discrete and continuous reservoir regimes [45, 46].
Collective Modes: Symmetric–Antisymmetric Decomposition and Interference.—Importantly, the total number of excitations in the system is conserved. As a result, the dynamics can be restricted to the single-excitation subspace, allowing for an exact and significantly simplified treatment of the energy exchange between the charger, the battery, and the structured environment. The total state of the system within the single-excitation sector can be written as
$|\Psi(t)\rangle$ $=$ $\alpha_{1}(t)\,|eg,\{0\}\rangle+\alpha_{2}(t)\,|ge,\{0\}\rangle$ (11)
$+$ $\sum_{k}\beta_{k}(t)\,|gg\rangle\otimes|1_{k}\rangle,$
where $|e\rangle_{j}$ and $|g\rangle_{j}$ denote the excited and ground states of the $j$-th two-level system, respectively. The states $|eg\rangle$ and $|ge\rangle$ represent configurations in which the excitation resides in the charger or in the battery, while the other qubit remains in its ground state. The state $|1_{k}\rangle$ corresponds to a single excitation in the $k$-th environmental mode, with both qubits in the ground state. This decomposition explicitly captures the energy exchange pathways between the charger, the battery, and the structured reservoir, and provides the basis for an exact treatment of the dynamics within the single-excitation manifold.
To facilitate the analysis of the system dynamics, it is advantageous to introduce the collective symmetric and antisymmetric basis. In this representation, the collective amplitudes are defined as $s(t)=(\alpha_{1}(t)+\alpha_{2}(t))/\sqrt{2}$ and $a(t)=(\alpha_{1}(t)-\alpha_{2}(t))/\sqrt{2}$ which correspond to the symmetric $|S\rangle=(|eg\rangle+|ge\rangle)/\sqrt{2}$ and antisymmetric $|A\rangle=(|eg\rangle-|ge\rangle)/\sqrt{2}$ collective states. This collective representation diagonalizes the coherent exchange interaction between the charger and the battery, thereby simplifying the mathematical structure of the problem. Within this basis, and under the near-resonant approximation, the system–environment interaction Hamiltonian can be compactly expressed as
$H_{SB}=\sum_{k}\sqrt{2}\,g_{k}\left[\cos(k_{0}d)\,\sigma_{S}^{+}+i\sin(k_{0}d)\,\sigma_{A}^{+}\right]b_{k}+\mathrm{H.c.},$ (12)
@@
where the collective raising operators are defined as $\sigma_{S,A}^{+}=(\sigma_{1}^{+}\pm\sigma_{2}^{+})/\sqrt{2}$. This form of the interaction Hamiltonian explicitly reveals how the geometric dependence enters the dynamics through the factors $\cos(k_{0}d)$ and $\sin(k_{0}d)$, which determine the effective coupling strength of the symmetric and antisymmetric collective channels to the environment. This expression shows that the symmetric and antisymmetric collective channels interact independently with the reservoir, and that the coupling strength of each channel depends explicitly on the interqubit separation. The geometric dependence enters the dynamics through the factors $\cos(k_{0}d)$ and $\sin(k_{0}d)$, such that the symmetric channel couples to the environment with amplitude proportional to $\cos(k_{0}d)$, while the antisymmetric channel couples with amplitude proportional to $\sin(k_{0}d)$.
For specific separations $d=n\lambda_{0}/2$, where $\sin(k_{0}d)=0$, the antisymmetric channel becomes completely decoupled from the reservoir and thus forms a geometry-protected dark state, whereas the symmetric channel remains coupled to the bath. Conversely, for separations $d=(2n+1)\lambda_{0}/4$, where $\cos(k_{0}d)=0$, the symmetric channel becomes fully decoupled and forms the dark state, while the antisymmetric channel continues to interact with the environment.
@@
The full details of the dynamical analysis are presented in the Appendix.A, where the derivation of the equations of motion, underlying assumptions, and intermediate analytical steps are provided in a systematic and comprehensive manner. Taking the partial trace over the environmental and charger degrees of freedom in Eq. (11), the reduced density matrix describing the quantum battery is obtained as follows
$\rho_{B}(t)=\begin{pmatrix}P_{B}(t)&0\\ 0&1-P_{B}(t)\end{pmatrix}$, (13)
@@
where $P_{B}(t)=|\alpha_{2}(t)|^{2}$ denotes the excitation probability of the quantum battery. The internal energy of the quantum battery is then given by
$E_{B}(t)=\omega_{0}|\alpha_{2}(t)|^{2}.$ (14)
@@
$\mathcal{W}(t)=(2|\alpha_{2}(t)|^{2}-1)\Theta(|\alpha_{2}(t)|^{2}-\frac{1}{2}),$ (15)
@@
$\mathcal{P}_{B}(t)=2\omega_{0}\,\Re\!\big[\alpha_{2}^{*}(t)\,\dot{\alpha}_{2}(t)\big].$ (16)
@@
$P^{\max}=\max_{\tau}\left[P(\tau)\right],$ (18)
which determines the highest achievable rate of energy transfer during the charging process.
For optimal operation of the quantum battery, both $\Delta E_{B}^{\max}$ and $\mathcal{W}^{\max}$ should attain large values, ensuring substantial energy storage and maximal extractable work, while a large $P^{\max}$ guarantees rapid charging. Together, these quantities provide a comprehensive evaluation of the battery’s energetic and dynamical performance.
In Fig. 2(a), we present the time-geometry plot of the internal energy variation of the quantum battery, $\Delta E_{B}(t)$, for the parameters $g=0.05\,\omega_{0}$, $J=-0.005\,\omega_{0}$, and $\lambda/g=0.02$. The small value of $\lambda$ places the system in the weak-dissipation, non-Markovian regime, where reservoir memory effects significantly influence the dynamics. The figure reveals a pronounced modulation of the charging process as a function of the geometric parameter $d/\lambda_{0}$. The vertical band structure indicates that the efficiency of energy transfer from the charger to the battery is strongly governed by the relative spatial phase $k_{0}d$. This behavior originates from the geometry-dependent collective decay rates, $\Gamma_{S}\propto\cos^{2}(k_{0}d)$ and $\Gamma_{A}\propto\sin^{2}(k_{0}d)$, which determine the effective coupling of the symmetric and antisymmetric collective channels to the structured reservoir. Consequently, constructive interference enhances energy transfer at specific separations, while destructive interference suppresses it, giving rise to geometry-induced dark configurations in which charging is strongly inhibited. The oscillatory structures along the temporal axis reflect coherent energy exchange between the charger and the battery that persists due to the long reservoir correlation time ($\lambda\ll g$). The appearance of revival-like features signals partial information and energy backflow from the reservoir, which is a characteristic signature of non-Markovian dynamics. Notably, both the maximum stored energy and the optimal charging time depend sensitively on the inter-qubit separation. Overall, the results demonstrate that in a structured environment the spatial geometry of the system acts as an effective control parameter for tuning the charging efficiency, dynamical stability, and the interplay between coherent exchange and dissipation. In particular, optimal charging occurs at separations where collective bright-state coupling is maximized, whereas dark-state configurations suppress energy transfer despite finite system–environment interaction.
Figure 2: Time-geometry plot of (a) the internal energy change $\Delta E_{B}(t)$ and (b) the ergotropy $\mathcal{W}(t)$. The battery is initially in its ground state and the charger excited. The parameters are $\omega_{0}=1$, $g=0.05\,\omega_{0}$, $J=-0.005\,\omega_{0}$, and $\lambda/g=0.02$.
@@
Figure 3: (a) the internal energy change $\Delta E_{B}(t)$ and (b) the ergotropy $\mathcal{W}(t)$, as function of $\lambda t$ for different values of geometry parameter. The battery is initially in its ground state and the charger excited. The parameters are $g=0.05\,\omega_{0}$, $J=-0.005\,\omega_{0}$, and $\lambda/g=0.02$.
Fig. 3(a) shows the internal energy change of the quantum battery, $\Delta E_{B}(t)$, for three representative geometric configurations defined by $\sin(k_{0}d)=0$, $\cos(k_{0}d)=0$, and $\sin(2k_{0}d)=1$. The parameters $\omega_{0}=1$, $g=0.05\,\omega_{0}$, $J=-0.005\,\omega_{0}$, and $\lambda/g=0.02$ correspond to a weak-dissipation, strongly non-Markovian regime, with the initial state $|eg\rangle$. The purely symmetric ($\sin(k_{0}d)=0$) and purely antisymmetric ($\cos(k_{0}d)=0$) configurations yield nearly identical charging dynamics. In both cases, one collective channel is completely dark, while the other remains fully coupled to the structured reservoir. Since the initial state contains equal symmetric and antisymmetric components, eliminating either channel effectively reduces the dynamics to a single active collective mode with comparable effective coupling strength. As a result, the internal energy evolution is symmetric under exchange of the two channels, leading to identical charging behavior. In contrast, the mixed configuration defined by $\sin(2k_{0}d)=1$ allows both collective channels to couple simultaneously and with equal strength. In this regime no dark-state protection occurs, and the symmetric and antisymmetric contributions jointly enhance the energy exchange between charger and battery. This results in larger oscillation amplitudes and improved charging performance. The long-lived oscillations observed in all cases arise from the small reservoir spectral width ($\lambda/g\ll 1$), which induces strong memory effects and enables repeated energy exchange between the system and the structured environment. Overall, the figure demonstrates that maximal charging is achieved when both collective channels contribute coherently, whereas single-channel (bright-only) configurations yield reduced performance despite full coupling strength.

### v2

# Spatial Phase Control of Energy and Ergotropy in Quantum Batteries

@@ Abstract
I Introduction
In recent years the thermodynamic effects and consequences arising from quantum evolution have emerged as a fundamental framework for the development of novel architectures for high-efficiency energy storage devices. These advances also provide the basis for the formulation and optimization of precise protocols for the control, conversion, and management of work and heat, approaches that open perspectives beyond the intrinsic limitations of classical thermodynamic processes [1, 2, 3, 4].
@@ I Introduction
Consequently, they have been extensively investigated and systematically analyzed in the contemporary literature [4, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Quantum batteries have been theoretically and experimentally investigated across a broad range of physical realizations. Proposed schemes include spin-chain models [21, 22, 23, 24] , atomic ensembles coupled to optical cavities [25, 26, 27, 28, 29] , and superconducting systems [30, 31, 32] . These diverse approaches rely on different interaction mechanisms and control techniques, thereby enabling the exploration of multiple strategies for microscopic energy storage and controlled energy transfer within quantum regimes.
@@
In this work, we investigate the role of the relative geometry between the charger and the quantum battery in a structured environment. In particular, we examine how the spatial seperation between the two qubits and the associated geometric phase influence the energy transfer process, charging dynamics and the amount of the extractable work of the battery through their control over collective interference and the spectral structure of the environment. Within this framework, we demonstrate that the system geometry not only regulates the effective coupling strength between the charger and the environment, but also, through the engineering of bright and dark collective channels, can suppress or enhance decay processes and energy backflow. As a result, the thermodynamic performance of the battery, including charging time, charging power, and maximum ergotropy, becomes directly governed by the geometric distance parameter. Furthermore, our analysis reveals that in the presence of environmental memory (non-Markovian effects), the system geometry can give rise to distinct regions in parameter space where the revival of extractable work occurs. In this sense, geometry emerges as a fundamental control resource, enabling the active engineering and optimization of quantum battery performance.
II Figures of merit
In this section, we present the thermodynamic framework employed to analyze quantum batteries. The system under consideration is a quantum battery characterized by the Hamiltonian $H_{B}$, with its state described by the density operator $\rho_{B}$. This quantum state can be expressed through its spectral decomposition in terms of its eigenvalues $r_{j}$ and corresponding eigenvectors $|r_{j}\rangle$
@@ II Figures of merit
III Physical model
Let us consider a composite quantum system consisting of two identical qubits located at $x_{1}=-l$ and $x_{2}=+l$, respectively. Within the quantum battery paradigm, one qubit is designated as the charger, while the other will be considered as the quantum battery. This physical setup is equivalent to two qubits placed inside a mirror-terminated one-dimensional waveguide of length $L$ [45]. We set $\hbar=1$ and work within the dipole and rotating-wave approximations. The total Hamiltonian of the system can be written as
@@ III Physical model
Figure 1: Schematic of the physical model: two identical qubits systems separated by a distance $2l$ inside a mirror-terminated one-dimensional waveguide. Qubit 1 acts as the charger (initially excited) and qubit 2 as the quantum battery (initially in the ground state).
where $H_{S}$ describes the free Hamiltonian of the two qubits (charger and battery), $H_{B}$ represents the discrete bosonic modes of the structured environment, and $H_{I}$ accounts for the interaction between the qubits and the environmental modes. The system Hamiltonian reads [45]
$H_{S}=\omega_{0}(\sigma_{1}^{+}\sigma_{1}^{-}+\sigma_{2}^{+}\sigma_{2}^{-})+\zeta(\sigma_{1}^{+}\sigma_{2}^{-}+\sigma_{1}^{-}\sigma_{2}^{+}),$ (7)
where $\omega_{0}$ is the transition frequency of the two-level systems. The parameter $\zeta$ represents the coherent dipole-dipole interaction [46]. In the quantum battery interpretation adopted here, qubit $1$ acts as the charger, while qubit $2$ represents the battery. The operators $\sigma_{j}^{\pm}$ denote the raising and lowering operators of the $j$-th qubit. The bath Hamiltonian is given by
@@
$H_{SB}=\sum_{k}\left[\xi_{k}^{(1)}\sigma_{1}^{-}a_{k}^{\dagger}+\xi_{k}^{(2)}\sigma_{2}^{-}a_{k}^{\dagger}+h.c.\right].$ (9)
@@
The model is schematically illustrated in Fig. 1. The Effects of geometry is encoded in the coupling amplitudes $\xi_{k}^{(j)}=\xi_{k}e^{ikx_{j}}$, with $x_{j}$ showing the location of the $j$-th qubit, where $k$ is the wave number [47, 48, 49, 50]. Since the interaction of the qubits predominantly occurs with modes near their transition frequency $\omega_{0}$, we simplify the analytical treatment by approximating the wave number in the position-dependent phase terms as $k\approx k_{0}=2\pi/\lambda_{0}$. Here, $\lambda_{0}$ denotes the resonant wavelength associated with the qubit transition frequency. Consequently, the position-dependent coupling coefficients can be expressed as [51, 52]
$\xi_{k}^{(1)}\simeq\xi_{k}e^{+ik_{0}l},\quad\xi_{k}^{(2)}\simeq\xi_{k}e^{-ik_{0}l}.$ (10)
To simplify the analysis of the system dynamics, it is convenient to introduce the collective symmetric and antisymmetric basis states, denoted by $|+\rangle=(|e_{c}g_{B}\rangle+|g_{c}e_{B}\rangle)/\sqrt{2}$ and $|-\rangle=(|e_{c}g_{B}\rangle-|g_{c}e_{B}\rangle)/\sqrt{2}$, respectively. In this basis, the dynamics can be decomposed into independent channels corresponding to symmetric and antisymmetric contributions, which significantly simplifies the mathematical treatment of the problem. Moreover, this transformation allows for a clearer physical interpretation of the energy transfer process and the role of coherent and dissipative interactions in the charging dynamics of the quantum battery. Within this basis, and under the near-resonant approximation, the system–environment interaction Hamiltonian can be compactly expressed as
$H_{SB}=\sum_{k}\sqrt{2}\,\left[\Gamma_{s}\,\Xi_{s}^{+}+\Gamma_{a}\,\Xi_{a}^{+}\right]a_{k}+\mathrm{H.c.},$ (11)
@@
where $\Gamma_{s}=\xi_{k}\cos(k_{0}l)$, $\Gamma_{a}=i\xi_{k}\sin(k_{0}l)$ and the collective raising operators are defined as $\Xi_{s}^{+}=(\sigma_{1}^{+}+\sigma_{2}^{+})/\sqrt{2}$ and $\Xi_{a}^{+}=(\sigma_{1}^{+}-\sigma_{2}^{+})/\sqrt{2}$. This form of the interaction Hamiltonian explicitly reveals how the geometric dependence enters the dynamics through the factors $\cos(k_{0}l)$ and $\sin(k_{0}l)$, which determine the effective coupling strength of the symmetric and antisymmetric collective channels to the environment. This expression shows that the symmetric and antisymmetric collective channels interact independently with the reservoir, and that the coupling strength of each channel depends explicitly on the interqubit separation. In this framework, we investigate the dynamical evolution of the composite system by restricting the analysis to the single-excitation sector, while assuming that the reservoir is initially in its vacuum state. Accordingly, the total initial state of the combined system and environment can be expressed as
$|\psi(0)\rangle=\left[c_{1}(0)\,|g_{B}\rangle|e_{c}\rangle+c_{2}(0)\,|e_{B}\rangle|g_{c}\rangle\right]\otimes|0_{k}\rangle_{E},$ (12)
where $|g_{c}\rangle$ ($|g_{B}\rangle$) and $|e_{c}\rangle$ ($|e_{B}\rangle$) denote the vacuum and single-excitation states of the charger (Battery), respectively. The state $|0_{k}\rangle_{E}$ corresponds to the vacuum state of the environment. The coefficients $c_{1}(0)$ and $c_{2}(0)$ represent the corresponding probability amplitudes. Within the single-excitation subspace, the time-dependent state of the total system can be expressed as
$|\psi(t)\rangle$ $=$ $\left[c_{1}(t)\,|g_{B},e_{c}\rangle+c_{2}(t)\,|e_{B},g_{c}\rangle\right]\otimes|0_{k}\rangle_{E}$ (13)
$+$ $\sum_{k}c_{k}(t)\,|g_{B},g_{c}\rangle\otimes|1_{k}\rangle_{E},$
where $|1_{k}\rangle_{E}$ denotes the state of the bosonic environment in which a single excitation occupies the $k$th mode. The coefficients $c_{1}(t)$, $c_{2}(t)$, and $c_{k}(t)$ represent the corresponding time-dependent probability amplitudes. By substituting Eq. (13) into the Schrödinger equation and using the Hamiltonian defined in Eq. (9), the coupled equations governing the time evolution of the probability amplitudes are obtained as follows
$\dot{c}_{1}(t)$ $=-i\omega_{0}c_{1}(t)-i\zeta c_{2}(t)-i\sum_{k}\xi_{k}^{(1)}\,c_{k}(t),$ (14)
$\dot{c}_{2}(t)$ $=-i\omega_{0}c_{2}(t)-i\zeta c_{1}(t)-i\sum_{k}\xi_{k}^{(2)}\,c_{k}(t),$ (15)
$\dot{c}_{k}(t)$ $=-i\omega_{k}c_{k}(t)-i\Big[\xi_{k}^{(1)*}c_{1}(t)+\xi_{k}^{(2)*}c_{2}(t)\Big].$ (16)
The environment is assumed to be characterized by a Lorentzian spectral density given by
$J(\omega)=\frac{\gamma}{2\pi}\frac{\lambda^{2}}{(\omega_{0}-\omega)^{2}+\lambda^{2}},$ (17)
where $\gamma$ denotes the effective system–reservoir coupling strength, and $\lambda$ represents the spectral width. The parameter $\lambda$ is inversely related to the environmental memory time $\tau_{E}$, such that $\tau_{E}=\lambda^{-1}$. By applying the analytical approach introduced in Ref. [45], the coupled integro-differential equations governing the system dynamics can be solved. We consider the physically relevant initial configuration in which the battery is initially empty and prepared in its ground state, while the charger is fully excited and contains the entire initial energy of the composite system. Accordingly, the probability amplitudes satisfy the initial conditions
$c_{1}(0)=1,\qquad c_{2}(0)=0.$ (18)
This procedure allows one to obtain closed-form expressions for the time-dependent probability amplitudes $c_{1}(t)$ and $c_{2}(t)$, which describe the evolution of the excitation in the cavity mode and the quantum battery, respectively. These amplitudes fully characterize the dynamical behavior of the system within the single-excitation subspace and provide direct access to the relevant physical quantities, such as the stored energy, ergotropy, and charging power of the quantum battery. The analytical solutions also enable a detailed investigation of the role of system parameters and environmental effects, including the influence of non-Markovian memory and system-reservoir coupling strength, on the charging dynamics.
@@
Taking the partial trace over the environmental and charger degrees of freedom in Eq. (13), the reduced density matrix describing the quantum battery is obtained as follows
$\rho_{B}(t)=|c_{2}(t)|^{2}|e_{B}\rangle\langle e_{B}|+(1-|c_{2}(t)|^{2})|g_{B}\rangle\langle g_{B}|$ (19)
@@
where $|c_{2}(t)|^{2}$ denotes the excitation probability of the quantum battery. The internal energy of the quantum battery is then given by
$E_{B}(t)=\omega_{0}|c_{2}(t)|^{2}.$ (20)
@@
$\mathcal{W}(t)=(2|c_{2}(t)|^{2}-1)\Theta(|c_{2}(t)|^{2}-\frac{1}{2}),$ (21)
@@
$\mathcal{P}_{B}(t)=2\omega_{0}\,\Re\!\big[c_{2}^{*}(t)\,\dot{c}_{2}(t)\big].$ (22)
@@
$\mathcal{P}^{\max}=\max_{\tau}\left[\mathcal{P}(t)\right],$ (24)
which determines the highest achievable rate of energy transfer during the charging process. For optimal operation of the quantum battery, both $\Delta E_{B}^{\max}$ and $\mathcal{W}^{\max}$ should attain large values, ensuring substantial energy storage and maximal extractable work, while a large $P^{\max}$ guarantees rapid charging. Together, these quantities provide a comprehensive evaluation of the battery’s energetic and dynamical performance.
In Fig. 2(a), we present the time-geometry plot of the internal energy variation of the quantum battery, $\Delta E_{B}(t)$, for the parameters $\zeta=0.01\,\omega_{0}$ and $\lambda/\gamma=0.04$. The small value of $\lambda$ places the system in the strong coupling, non-Markovian regime, where reservoir memory effects significantly influence the dynamics. The figure reveals a pronounced modulation of the charging process as a function of the geometric parameter $l/\lambda_{0}$. The vertical band structure indicates that the efficiency of energy transfer from the charger to the battery is strongly governed by the relative spatial phase $k_{0}l$. This behavior originates from the geometry-dependent collective decay rates, $\Gamma_{s}\propto\cos(k_{0}l)$ and $\Gamma_{a}\propto\sin(k_{0}l)$, which determine the effective coupling of the symmetric and antisymmetric collective channels to the structured reservoir. Consequently, constructive interference enhances energy transfer at specific separations, while destructive interference suppresses it, giving rise to geometry-induced dark configurations in which charging is strongly inhibited. The oscillatory structures along the temporal axis reflect coherent energy exchange between the charger and the battery that persists due to the long reservoir correlation time ($\lambda\ll\gamma$). The appearance of revival-like features signals partial information and energy backflow from the reservoir, which is a characteristic signature of non-Markovian dynamics. Notably, both the maximum stored energy and the optimal charging time depend sensitively on the inter-qubit separation. Overall, the results demonstrate that in a structured environment the spatial geometry of the system acts as an effective control parameter for tuning the charging efficiency, dynamical stability, and the interplay between coherent exchange and dissipation. In particular, optimal charging occurs at separations where collective bright-state coupling is maximized, whereas dark-state configurations suppress energy transfer despite finite system–environment interaction.
Figure 2: Time-geometry plot of (a) the internal energy change $\Delta E_{B}(t)$ and (b) the ergotropy $\mathcal{W}(t)$. The battery is initially in its ground state and the charger excited. The parameters are $\zeta=0.01\,\omega_{0}$ and $\lambda/\gamma=0.04$.
@@
Figure 3: (a) the internal energy change $\Delta E_{B}(t)$ and (b) the ergotropy $\mathcal{W}(t)$, as function of $\lambda t$ for different values of geometry parameter. The battery is initially in its ground state and the charger excited. The parameters are $\zeta=0.01\,\omega_{0}$, and $\lambda/\gamma=0.04$.
Fig. 3(a) shows the internal energy change of the quantum battery, $\Delta E_{B}(t)$, for three representative geometric configurations defined by $\Gamma_{a}=0$, $\Gamma_{s}=0$, and $\Gamma_{s}=-i\Gamma_{a}$. The parameters $\zeta=0.01\,\omega_{0}$, and $\lambda/\gamma=0.04$ correspond to a strong coupling, non-Markovian regime. The purely symmetric ($\Gamma_{a}=0$) and purely antisymmetric ($\Gamma_{s}=0$) configurations yield nearly identical charging dynamics. In both cases, one collective channel is completely dark, while the other remains fully coupled to the structured reservoir. Since the initial state contains equal symmetric and antisymmetric components, eliminating either channel effectively reduces the dynamics to a single active collective mode with comparable effective coupling strength. As a result, the internal energy evolution is symmetric under exchange of the two channels, leading to identical charging behavior. In contrast, the mixed configuration defined by $\Gamma_{s}=-i\Gamma_{a}$ allows both collective channels to couple simultaneously and with equal strength. In this regime no dark-state protection occurs, and the symmetric and antisymmetric contributions jointly enhance the energy exchange between charger and battery. This results in larger oscillation amplitudes and improved charging performance. The long-lived oscillations observed in all cases arise from the small reservoir spectral width ($\lambda/\gamma\ll 1$), which induces strong memory effects and enables repeated energy exchange between the system and the structured environment. Overall, the figure demonstrates that maximal charging is achieved when both collective channels contribute coherently, whereas single-channel (bright-only) configurations yield reduced performance despite full coupling strength.